OpenStudy (anonymous):

If the cos 30° is 0.866, then the sin _________ ° is 0.866.

OpenStudy (anonymous):

@sidsiddhartha @ganeshie8 @Rohith15

OpenStudy (anonymous):

cos A = sin(90 - A) A = 30 Therefore, 90 - A = 60

OpenStudy (anonymous):

So i put 60 into the blank

OpenStudy (anonymous):

yeah..!!

OpenStudy (anonymous):

thnxs